1g metal combines with 0.67g of o2 determine the mass of metal carbonate produced by 1 g metal oxide

Grade:12th pass

2 Answers

Vikas TU
4 years ago
1 g of metal oxide gives 0.68 g of the metal
Hence Mass of Oxygen = 1- 0.68 = 0.32g
therefore 68g of metal combine with 32g of oxygen in the metal oxide.
Equivalent weight of the metal = (weight of metal / weight of oxygen combined )×8
Hence equivalent mass of the metal = (68/32)×8=17
merrin tressacherian
4 years ago
Equivalent weight of metal=(1/0.67)*8=11.94g
Mass of metal carbonate=(equivalent mass of metal carbonate/equivalent mass of metal oxide)*weight of the metal oxide=(42/20)*1=2.1